OpenStudy (anonymous):

Prove or Disprove. For any integer a, gcd(22a +7, 3a+1) = 1.

OpenStudy (anonymous):

22a+7=7*(3a+1)+a 3a+1=3*a+1 a=1*a+0 THUS, gcd(22a+7,3a+1)=1

OpenStudy (anonymous):

Thanks, can you also help me out with this? Prove or Disprove. For any integer a, gcd(a^2, a+1)=1

OpenStudy (anonymous):

First PLZ try yourself

OpenStudy (anonymous):

looool you should see the amount of time ive spent and paper

OpenStudy (anonymous):

i only do this as a last resort

OpenStudy (anonymous):

a^2 = (a-1)*(a+1)+1 a+1=1*(a+1)+0 Thus, gcd(a^2,a+1)=1

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

@Autrix ???